Two small identical conducting spheres are placed with their centers 0.50 m apart. One is given a charge of 12 ✕ 10−9 C, the other a charge of −20 ✕ 10−9 C.

(a) Find the electrostatic force exerted on one sphere by the other.

magnitude N
direction ---Select--- attractive repulsive


(b) The spheres are connected by a conducting wire. Find the electrostatic force between the two after equilibrium is reached, where both spheres have the same charge.

magnitude N
direction ---Select--- attractive repulsive
Answer #1

1 12 10C

220 10-C

d=0.5m

(a) kq142

F 91012 102010C/(0.5m)2

F-8.64 10N

Since the charge are unlike ,Force is Attractive

==========================

(b)After spheres reaches equilibrium ,charges change

Y1 = q2 = (12 * 10-9C-20 * 10-9C)/2 =-4 * 10-9c

kq142

F 9 * 109 * 4 * 10-C * 4 * 10-9C/(0.5m)2

F 5.76 10N

Since the charge are like(same),Force is Repulsive.

======================================